OpenStudy (anonymous):

Simplify your answer contain only positive exponents 3x^3 times 4x^-3

OpenStudy (anonymous):

\[x^a * x^b = x^{a+b}\] \[3x^3 * 4x^{-3} = 12x^{3-3} = 12x^0 = 12(1) = 12 = explicit\] this can also be seen as: also:\[3x^3 *\frac{ 4 }{ x^3 } = \frac{ 3*4x^3 }{ x^3 } = 12\] x^3 cancels out